Fraw Company owns 15% of the common stock of Trace Corporation and used the cost method to account for this investment. Trace reported net income of $110,000 for 2018 and paid dividends of $60,000 on October 1, 2018. How much income should Fraw recognize on this investment in 2018?

$16,500

$9,000

$50,000

$25,500.
Answer #1

Answer:- $7500

Explanation:-

Net income = 110000×15% = $16500

Cashs dividends = 60000 × 15% = $9000

INCOME FROM INVESTMENT = $ 16500 - 9000 = $ 7500